# The role of charm and unflavored mesons in prompt atmospheric lepton fluxes

### v1

@@ 2.3 Comparisons with IceCube data and upper limits
Figure 3 shows the angle-averaged atmospheric conventional and prompt $\mu^{+}+\mu^{-}$ flux (left panel) and $\nu_{\mu}+\bar{\nu}_{\mu}$ flux (right panel). For the conventional component, we average over zenith angles $\theta_{\rm zen}=0^{\circ}-60^{\circ}$ to avoid nearly horizontal directions where experimental mis-reconstruction of muon events may be significant IceCube:2015wro . In the left panel, the conventional flux, the pQCD(cc¯)($\bar{c}$) contribution, and the unflavored meson contribution are shown separately, along with the total flux (orange curve). The Sibyll-2.3c hadronic interaction model predicts roughly comparable contributions to the prompt atmospheric $\mu^{+}+\mu^{-}$ flux from heavy flavor and light unflavored mesons which makes them difficult to distinguish in the figure. The IceCube measurement of the angle-averaged muon flux for $\theta_{\rm zen}<60^{\circ}$ is shown by the data points with bin-by-bin errorbars IceCube:2015wro . Above $\sim 10^{5}$ GeV, the MCEq prediction underestimates the observed flux by a factor of $\sim 1.5-2$ (see the red points in fig. 6), although the statistical uncertainties are larger as the energy increases. The right panel of fig. 3 shows IceCube’s upper bound on the prompt neutrino flux assuming isotropy PhysRevD.78.043005 , a reliable approximation in the energy range shown in the figure. The IceCube upper limit on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux still allows room for additional prompt contributions. In what follows, we explore possible additional sources of prompt leptons, i.e. both heavy flavored hadrons generated by mechanisms involving intrinsic charm and light unflavored mesons, and study their interplay in the production of prompt atmospheric muons and muon neutrinos.
@@ 3 Intrinsic charm
Constraints on intrinsic charm contributions to the atmospheric neutrino flux have been discussed in the literature (see, e.g., refs. Ostapchenko:2022thy ; Garzelli:2023jlq ; Laha:2016dri ; Halzen:2016thi ). The potential discrepancy between the atmospheric muon measurements and the predictions from MCEq suggest missing contributions to the atmospheric muon flux. One approach is to add intrinsic charm contributions to the prompt atmospheric muon flux, adjusting the normalization of the intrinsic charm contributions to account for the deficit in the theory predictions for muons, then checking for consistency with the experimental upper bound on the flux of atmospheric muon neutrinos.
We consider a phenomenological implementation Garzelli:2023jlq ; Ostapchenko:2022thy of intrinsic charm in which valence-like momentum fractions are transferred to $\bar{D}^{0}$ and $\Lambda_{c}$ in proton-air ($pA$) scattering and to $D^{-}$ and $\Lambda_{c}$ in neutron-air ($nA$) scattering:
@@
Following ref. Ostapchenko:2022thy ; Garzelli:2023jlq , we write the differential cross section for proton-air production of charm hadron $h_{c}$ via non-perturbative intrinsic charm as
@@
Figure 4: $f_{h_{c}^{(intr}}$ of $D$ mesons and $\Lambda_{c}$ as a function of $x_{h_{c}}$ (see eq. 11) using the Regge ansatz Kaidalov:1985jg ; Kaidalov:2003wp with either $a_{N}=-0.5$ (Regge1) or $a_{N}=0$ (Regge2), and using the Hobbs, Londergan and Melnitchouk (HLM) approach Hobbs:2013bia (see text for details).
Several theoretical approaches exist to model the intrinsic charm fragmentation, e.g., Brodsky et al. (BHPS) Brodsky:1980pb ; Brodsky:1981se , Regge ansatz Kaidalov:1985jg ; Kaidalov:2003wp and meson-baryon models, e.g., as in ref. Hobbs:2013bia . We use the Regge ansatz as our starting point, and also show results for the meson-baryon model approach of Hobbs, Londergan and Melnitchouk (HLM) Hobbs:2013bia . The Regge fragmentation function (normalized to unity) is
@@
Hobbs, Londergan and Melnitchouk Hobbs:2013bia determined splitting functions in a meson-baryon model, for which the $\bar{D}^{*0}\Lambda_{c}$ dominates for $pA$ interactions. In fig. 4 we show the $\Lambda_{c}$ distribution using $x_{\Lambda_{c}}=1-x_{\bar{D}^{*0}}$. We account for a diminished $\bar{D}^{0}$ momentum fraction from the decay $\bar{D}^{*0}\to D^{0}\pi^{0}$ and $\bar{D}^{*0}\to D^{0}\gamma$ following refs. Cacciari:2003zu ; Cacciari:2005uk . For $nA$ interactions, $D^{*-}\Lambda_{c}$ fluctuations dominate. The decays of $\bar{D}^{*-}\to D^{-}\pi^{0}$, $\bar{D}^{*-}\to D^{-}\gamma$ and $\bar{D}^{*-}\to\bar{D}^{0}\pi^{-}$ have very similar kinematics and yield $D$ meson $x$-distributions that follows the HLM solid line in fig. 4. For both the Regge and HLM approaches, the charmed baryon carries a larger energy fraction than the charmed meson, although the energy balance is much closer in the HLM approach than with the Regge ansatz.
@@ 4.1 Angle-averaged atmospheric lepton fluxes
As we saw in fig. 3, the MCEq prediction for the atmospheric muon flux lies below the IceCube data. In this section, we begin by introducing intrinsic charm (IC) contributions to both muon and muon neutrino fluxes. The left panel of fig. 5 shows the best-fit intrinsic charm contribution to the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric muon flux, as well as the conventional, prompt light unflavored, prompt pQCD($c\bar{c})$ and the total muon flux contributions calculated using MCEq, along with the IceCube muon data IceCube:2015wro . The best-fit value of the $w_{c}^{\rm intr}$ parameter amounts to $w_{c}^{\rm intr}=3.87\times 10^{-3}$ for the default Regge1 intrinsic charm fragmentation model. This value for $w_{c}^{\rm intr}$ implies that $\sim 0.4\%$ of the nucleon-air cross section is due to intrinsic charm. Despite the small value of $w_{c}^{\rm intr}$, the energy distribution of the charm hadrons following eq. 11 is large because the cosmic ray flux falls so rapidly with increasing cosmic ray energy. Figure 6 shows the ratio of the IceCube data (including experimental error bars) to the MCEq prediction with the addition of intrinsic charm with $w_{c}^{\rm intr}=3.87\times 10^{-3}$.
@@
The $1\sigma$ interval for $w_{c}^{\rm intr}$ based on a $\chi^{2}$ fit is $(2.89-4.85)\times 10^{-3}$. Table 1 shows the best fit values and $1\sigma$ intervals for the Regge1, Regge2 and HLM models. All considered models lead to similar ranges of a few $\times 10^{-3}$ for $w_{c}^{\rm intr}$.
@@
Figure 5: As in fig. 3, the angle-averaged atmospheric fluxes with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Including intrinsic charm (IC) contribution with the weight $w_{c}^{\rm intr}=3.87\times 10^{-3}$, the best-fit for the IceCube data IceCube:2015wro . The gray curve shows the sum of conventional and prompt $\mu^{+}+\mu^{-}$ fluxes including intrinsic charm. Right: The prompt atmospheric muon neutrino flux including intrinsic charm contribution with $w^{c}_{\rm intr}=3.87\times 10^{-3}$ and the upper limit on the prompt flux from IceCube IceCube:2016umi . The sum of perturbative and intrinsic charm contributions to the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux is shown with the gray curve.
@@
On the other hand, the right panel of fig. 5 shows that the prompt perturbative plus intrinsic charm contributions to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux significantly overshoot the IceCube upper bound IceCube:2016umi . To satisfy the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux constraint from ref. IceCube:2016umi , $w_{c}^{\rm intr}$ must be reduced, which also implies that intrinsic charm alone cannot account for the apparent deficit in the flux of atmospheric muons at high energies. The IceCube upper bound IceCube:2015wro relies on the shape of the prompt flux from Enberg et al. Enberg:2008te as a function of neutrino energy. We observe that this flux has a slightly different shape than the MCEq prediction. We normalize $w_{c}^{\rm intr}$ to the value of the IceCube upper bound at $E_{\nu}=10^{4}$ GeV, where the upper bound has a minimum. Figure 7 shows the intrinsic charm contributions to the atmospheric muon neutrino (left) and muon (right) fluxes for $w_{c}^{\rm intr}=4.65\times 10^{-4}$. Regge2 and HLM fragmentation models have similar best-fit values to Regge1. With $w_{c}^{\rm intr}$ nearly a factor of 10 smaller than the $\mu^{+}+\mu^{-}$ best-fit value, the atmospheric muon flux data remain above the sum of the MCEq flux and this now reduced intrinsic charm contribution, as shown in the right panel of fig. 7. Figure 8 shows the ratio of IceCube muon data with error bars to the muon flux with and without intrinsic charm, as in fig. 6, but here with $w_{c}^{\rm intr}=4.65\times 10^{-4}$. Such a small value of $w_{c}^{\rm intr}$ has little impact on the muon flux ratio.
Table 1: Values of $w_{c}^{\rm intr}$ for intrinsic-charm model-dependent fit to the angle-averaged muon energy spectrum ($\theta_{\rm zen}\leq$ $60^{\circ}$). Best fit values of $w_{c}^{\rm intr}$ along with the $1\,\sigma$ interval are shown for the two Regge ansatz models Kaidalov:1985jg ; Kaidalov:2003wp and for the Hobbs, Londergan and Melnitchouk (HLM) model Hobbs:2013bia .
@@
HLM 2.99 2.233.742.23-3.74
Figure 6: Ratio of IceCube muon data with experimental error bars IceCube:2015wro to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=3.87\times 10^{-3}$.
Figure 7: As in fig. 3, angle-averaged with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Comparison of prompt neutrino spectrum ($\nu_{\mu}+\bar{\nu}_{\mu}$) including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$ with the upper limit from IceCube IceCube:2016umi . Right: Comparison of angle-averaged muon flux ($\mu^{+}+\mu^{-}$) with IceCube data IceCube:2015wro after including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$.
Figure 8: Ratio of IceCube muon data with experimental error bars IceCube:2015wro to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$.
@@
Figure 9: Left: The angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux ($\mu^{+}+\mu^{-}$) including an intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$ and scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=4.33$. Right: Without any intrinsic charm contribution, the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux with scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=4.80$.
[…]
Model $\mathbf{{\it w_{c}}^{\rm intr}\times 10^{-4}}$ $\mathbf{\alpha_{\rm unfl}}$ $\mathbf{1\sigma}$ interval

### v2

@@ 2.3 Comparisons with IceCube data and upper limits
Figure 3 shows the angle-averaged atmospheric conventional and prompt $\mu^{+}+\mu^{-}$ flux (left panel) and $\nu_{\mu}+\bar{\nu}_{\mu}$ flux (right panel). For the conventional component, we average over zenith angles $\theta_{\rm zen}=0^{\circ}-60^{\circ}$ to avoid nearly horizontal directions where experimental mis-reconstruction of muon events may be significant Aartsen and others (2016a). In the left panel, the conventional flux, the pQCD(cc¯)($\bar{c}$) contribution, and the unflavored meson contribution are shown separately, along with the total flux (orange curve). The Sibyll-2.3c hadronic interaction model predicts roughly comparable contributions to the prompt atmospheric $\mu^{+}+\mu^{-}$ flux from heavy flavor and light unflavored mesons which makes them difficult to distinguish in the figure. The IceCube measurement of the angle-averaged muon flux for $\theta_{\rm zen}<60^{\circ}$ is shown by the data points with bin-by-bin errorbars Aartsen and others (2016a). Above $\sim 10^{5}$ GeV, the MCEq prediction underestimates the observed flux by a factor of $\sim 1.5-2$ (see the red points in fig. 6), although the statistical uncertainties are larger as the energy increases. The right panel of fig. 3 shows IceCube’s 2025 upper bound on the prompt neutrino flux Abbasi and others (2025a) assuming isotropy Enberg et al. (2008), a reliable approximation in the energy range shown in the figure. The IceCube upper limit on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux still allows room for additional prompt contributions. In what follows, we explore possible additional sources of prompt leptons, i.e., both heavy flavored hadrons generated by mechanisms involving intrinsic charm and light unflavored mesons, and study their interplay in the production of prompt atmospheric muons and muon neutrinos.
@@ 3 Intrinsic charm
Constraints on intrinsic charm contributions to the atmospheric neutrino flux have been discussed in the literature (see, e.g., refs. Ostapchenko et al. (2023); Garzelli et al. (2024); Laha and Brodsky (2017); Halzen and Wille (2016)). The potential discrepancy between the atmospheric muon measurements and the predictions from MCEq suggest missing contributions to the atmospheric muon flux. One approach is to add intrinsic charm contributions to the prompt atmospheric muon flux, adjusting the normalization of the intrinsic charm contributions to account for the deficit in the theory predictions for muons, then checking for consistency with the experimental upper bound on the flux of atmospheric muon neutrinos.
We consider a phenomenological implementation Garzelli et al. (2024); Ostapchenko et al. (2023) of intrinsic charm in which valence-like momentum fractions are transferred to $\bar{D}^{0}$ and $\Lambda_{c}$ in proton-air ($pA$) scattering and to $D^{-}$ and $\Lambda_{c}$ in neutron-air ($nA$) scattering:
@@
Following ref. Ostapchenko et al. (2023); Garzelli et al. (2024), we write the differential cross section for proton-air production of charm hadron $h_{c}$ via non-perturbative intrinsic charm as
@@
Figure 4: $f_{h_{c}^{(intr}}$ of $D$ mesons and $\Lambda_{c}$ as a function of $x_{h_{c}}$ (see eq. 11) using the Regge ansatz Kaidalov and Piskunova (1986); Kaidalov (2003) with either $a_{N}=-0.5$ (Regge1) or $a_{N}=0$ (Regge2), and using the Hobbs, Londergan and Melnitchouk (HLM) approach Hobbs et al. (2014) (see text for details).
Several theoretical approaches exist to model the intrinsic charm fragmentation, e.g., Brodsky et al. (BHPS) Brodsky et al. (1980, 1981), Regge ansatz Kaidalov and Piskunova (1986); Kaidalov (2003) and meson-baryon models, e.g., as in ref. Hobbs et al. (2014). We use the Regge ansatz as our starting point, and also show results for the meson-baryon model approach of Hobbs, Londergan and Melnitchouk (HLM) Hobbs et al. (2014). The Regge fragmentation function (normalized to unity) is
@@
Hobbs, Londergan and Melnitchouk Hobbs et al. (2014) determined splitting functions in a meson-baryon model, for which the $\bar{D}^{*0}\Lambda_{c}$ dominates for $pA$ interactions. In fig. 4 we show the $\Lambda_{c}$ distribution using $x_{\Lambda_{c}}=1-x_{\bar{D}^{*0}}$. We account for a diminished $\bar{D}^{0}$ momentum fraction from the decay $\bar{D}^{*0}\to D^{0}\pi^{0}$ and $\bar{D}^{*0}\to D^{0}\gamma$ following refs. Cacciari and Nason (2003); Cacciari et al. (2006). For $nA$ interactions, $D^{*-}\Lambda_{c}$ fluctuations dominate. The decays of $\bar{D}^{*-}\to D^{-}\pi^{0}$, $\bar{D}^{*-}\to D^{-}\gamma$ and $\bar{D}^{*-}\to\bar{D}^{0}\pi^{-}$ have very similar kinematics and yield $D$ meson $x$-distributions that follows the HLM solid line in fig. 4. For both the Regge and HLM approaches, the charmed baryon carries a larger energy fraction than the charmed meson, although the energy balance is much closer in the HLM approach than with the Regge ansatz.
@@ 4.1 Angle-averaged atmospheric lepton fluxes
As we saw in fig. 3, the MCEq prediction for the atmospheric muon flux lies below the IceCube data. In this section, we begin by introducing intrinsic charm (IC) contributions to both muon and muon neutrino fluxes. The left panel of fig. 5 shows the best-fit intrinsic charm contribution to the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric muon flux, as well as the conventional, prompt light unflavored, prompt pQCD($c\bar{c})$ and the total muon flux contributions calculated using MCEq, along with the IceCube muon data Aartsen and others (2016a). The best-fit value of the $w_{c}^{\rm intr}$ parameter amounts to $w_{c}^{\rm intr}=3.93\times 10^{-3}$ for the default Regge1 intrinsic charm fragmentation model. This value for $w_{c}^{\rm intr}$ implies that $\sim 0.4\%$ of the nucleon-air cross section is due to intrinsic charm. Despite the small value of $w_{c}^{\rm intr}$, the energy distribution of the charm hadrons following eq. 11 is large because the cosmic ray flux falls rapidly with increasing cosmic ray energy. Figure 6 shows the ratio of the IceCube data (including experimental error bars) to the MCEq prediction both without and with the addition of an intrinsic charm contribution corresponding to $w_{c}^{\rm intr}=3.93\times 10^{-3}$.
@@
The $1\sigma$ interval for $w_{c}^{\rm intr}$ based on a $\chi^{2}$ fit is $(2.95-4.91)\times 10^{-3}$. Table 1 shows the best fit values and $1\sigma$ intervals for the Regge1, Regge2 and HLM models. All considered models lead to similar ranges of a few $\times 10^{-3}$ for $w_{c}^{\rm intr}$.
@@
Figure 5: As in fig. 3, the angle-averaged atmospheric fluxes with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Including intrinsic charm (IC) contribution with the weight $w_{c}^{\rm intr}=3.93\times 10^{-3}$, the best-fit for the IceCube muon data Aartsen and others (2016a). The gray curve shows the sum of conventional and prompt $\mu^{+}+\mu^{-}$ fluxes including intrinsic charm. Right: The prompt atmospheric muon neutrino flux including intrinsic charm contribution with $w_{c}^{\rm intr}=3.93\times 10^{-3}$ and the upper limit on the prompt flux from IceCube Abbasi and others (2025a). The sum of perturbative and intrinsic charm contributions to the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux is shown with the gray curve.
@@
On the other hand, the right panel of fig. 5 shows that the prompt perturbative plus intrinsic charm contributions to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux significantly overshoot the IceCube upper bound Abbasi and others (2025a). To satisfy the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux constraint from ref. Abbasi and others (2025a), $w_{c}^{\rm intr}$ must be reduced, which also implies that intrinsic charm alone cannot account for the apparent deficit in the flux of atmospheric muons at high energies. The IceCube upper bound Abbasi and others (2025a) shown in the right panel of fig. 5 relies on the shape of the prompt flux evaluated using charm production from Sibyll-2.3c. The best-fit intrinsic charm contribution that saturates the IceCube upper bound is $w_{c}^{\rm intr}=1.01\times 10^{-3}$. Figure 7 shows the intrinsic charm contributions to the atmospheric muon neutrino (left) and muon (right) fluxes for $w_{c}^{\rm intr}=1.01\times 10^{-3}$. The Regge2 and HLM fragmentation models have similar best-fit values to the Regge1 one. The value of $w_{c}^{\rm intr}$ is $\sim 1/4$ of the $\mu^{+}+\mu^{-}$ best-fit value, so the atmospheric muon flux data remain above the sum of the flux contributions, now with a reduced intrinsic charm component, as shown in the right panel of fig. 7.
As noted above, the IceCube limit on the prompt atmospheric neutrino flux in ref. Abbasi and others (2025a) is a scaling of the prompt flux evaluated with Sibyll-2.3c. On the other hand, the 2016 IceCube upper bound Aartsen and others (2016b) relied on the shape of the prompt flux from Enberg et al. (ERS) Enberg et al. (2008) as a function of neutrino energy. The slightly different shapes of the two prompt neutrino fluxes can be seen with the aforementioned upper bounds in fig. 8, where the solid red bound is from ref. Abbasi and others (2025a) while the dashed red bound is from ref. Aartsen and others (2016b). For the 2016 upper bound, we normalize $w_{c}^{\rm intr}$ to the value of the IceCube upper bound at $E_{\nu}=10^{4}$ GeV, where this bound has a minimum. The scaled intrinsic charm and pQCD($c\bar{c}$) plus intrinsic charm contributions to the prompt neutrino flux are shown in fig. 8 with dashed lines corresponding to $w_{c}^{\rm intr}=5.34\times 10^{-4}$ resulting from saturating this IceCube bound Aartsen and others (2016b). In the case of the 2025 bound, we normalize $w_{c}^{\rm intr}$ in such a way that the total prompt neutrino contribution (pQCD + intrinsic charm) saturates the bound at $E_{\nu}=10^{5}$ GeV, obtaining $w_{c}^{\rm intr}=1.01\times 10^{-3}$ and the intrinsic charm component shown by solid line in the same figure.
Figure 9 shows the ratio of IceCube muon data with error bars to the muon flux with and without intrinsic charm, as in fig. 6, but here with $w_{c}^{\rm intr}=1.01\times 10^{-3}$. Such a small value of $w_{c}^{\rm intr}$ has little impact on the muon flux ratio.
@@
HLM 3.03 2.283.792.28-3.79
Table 1: Values of $w_{c}^{\rm intr}$ for intrinsic-charm model-dependent fit to the angle-averaged muon energy spectrum ($\theta_{\rm zen}\leq$ $60^{\circ}$). Best fit values of $w_{c}^{\rm intr}$ along with the $1\,\sigma$ interval are shown for the two Regge ansatz models Kaidalov and Piskunova (1986); Kaidalov (2003) and for the Hobbs, Londergan and Melnitchouk (HLM) model Hobbs et al. (2014).
Figure 6: Ratio of IceCube muon data with experimental error bars Aartsen and others (2016a) to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=3.93\times 10^{-3}$.
Figure 7: As in fig. 3, angle-averaged with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Comparison of prompt neutrino spectrum ($\nu_{\mu}+\bar{\nu}_{\mu}$) including intrinsic charm contribution with $w_{c}^{\rm intr}=1.01\times 10^{-3}$ with the upper limit from IceCube Abbasi and others (2025a). Right: Comparison of angle-averaged muon flux ($\mu^{+}+\mu^{-}$) with IceCube data Aartsen and others (2016a) after including intrinsic charm contribution with $w_{c}^{\rm intr}=1.01\times 10^{-3}$.
Figure 8: Comparisons of prompt neutrino spectrum ($\nu_{\mu}+\bar{\nu}_{\mu}$) including intrinsic charm contribution with $w_{c}^{\rm intr}=1.01\times 10^{-3}$ to the 2025 upper limit from IceCube Abbasi and others (2025a), and with $w_{c}^{\rm intr}=5.34\times 10^{-4}$ to the 2016 upper bound from IceCube Aartsen and others (2016b) that relies on the ERS prompt flux spectrum Enberg et al. (2008).
Figure 9: Ratio of IceCube muon data with experimental error bars Aartsen and others (2016a) to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=1.01\times 10^{-3}$.
@@
Figure 10: Left: The angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux ($\mu^{+}+\mu^{-}$) including an intrinsic charm contribution with $w_{c}^{\rm intr}=1.01\times 10^{-3}$ and scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=3.84$. Right: Without any intrinsic charm contribution, the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux with scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=4.86$.
Model $\mathbf{{\it w_{c}}^{\rm intr}\times 10^{-3}}$ $\mathbf{\alpha_{\rm unfl}}$ $\mathbf{1\sigma}$ interval